# Implementation of full and simplified likelihoods in CheckMATE Iñaki Lara and Krzysztof Rolbiecki

### v2

@@ 3 Validation
In Figure 4, we show validation plots for slepton search regions. The electron and muon channels are shown separately. Very good agreement between ATLAS and CheckMATE is obtained for the shape-fit analysis. Additionally, by comparing red exclusion curves obtained using the best-SR-method, we find a clear benefit of using multibin signal regions. Since there is a negligible difference between expected and observed results (both for ATLAS and CheckMATE) we only show the observed limits.
In Figure 5, we show validation plots for the wino-like chargino-neutralino pair production (EW search region). The searches for different parities of LSP and NLSP are shown separately: upper row $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})>0$ (same parity) and $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})<0$ (opposite parity) in the lower row; see eg. Ref. [78] for a discussion of CP properties of the neutralino sector and consequences for the decay kinematics. The electron and muon channels are shown separately. These examples show a danger of using a simplified likelihood model, which clearly produces too strong constraints.
Figure 4: Validation plots (observed limits) for the search atlas_1911_12606 (SUSY-2018-16), slepton pair production. Left panel: smuon search; right panel selectron search.
Figure 5: Validation plots for the search atlas_1911_12606 (SUSY-2018-16). Top row: $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})>0$; bottom row: $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})<0$. Left panels: observed limits; right panels: expected limits.
@@
This is a search [35] for hadronically decaying supersymmetric partners of top quark, and up-type, 3rd generation scalar leptoquark. The final states consist of several jets, 0 leptons, and large $E_{\mathrm{T}}^{\mathrm{miss}}$. The requirements for $b$ jets vary between signal regions. There are 3 multibin signal regions: SRA-B which targets scenarios with highly boosted top quarks in the final state and is divided into 6 bins according to an invariant mass of a large-$R$ jet; SRC for scenarios with 3-body decays of stops which is divided into 5 bins according to a recursive jigsaw reconstruction technique variable $R_{\textrm{ISR}}$ [79]; SRD for scenarios with compressed spectra and 4-body decays of stops which is divided into 3 bins according to the number of identified $b$-jets. No likelihood model is provided, and only simplified likelihood fitting is available.
The validation plots for this search are shown in Figure 6. We include the full range of stop masses and decay modes. A good agreement across the parameter plane is observed. For the high-stop-mass region the CheckMATE result is slightly weaker than ATLAS, clearly extending the exclusion reach compared to single SR exclusion. In the more compressed scenarios, the simplified fit gives a slightly too strong exclusion in some regions of the parameter space, however, this is below 30% difference in the exclusion strength.
@@
This is a search [36] for supersymmetric partners of top quark decaying to a Higgs or $Z$ boson. The final state Higgs boson is reconstructed from a pair of $b$-jets while the $Z$ boson from a same-flavor opposite-sign dilepton pair. There are 3 multibin signal regions that are shape fits in $E_{\mathrm{T}}^{\mathrm{miss}}$, $E_{\mathrm{T}}^{\mathrm{miss}}$-significance and $p_{\mathrm{T}}$ of the $Z$ candidate. No likelihood model is provided, and only a simplified likelihood model is available.
@@
Figure 7: Validation plots for the search atlas_2006_05880 (SUSY-2018-31). Left panels: observed limits; right panels: expected limits.
@@
This is a search [37] for squarks (1st and 2nd generation) and gluinos in final states with 2-6 jets, 0 leptons and missing transverse momentum. There are three multibin signal regions in this search: MB-SSd which targets production of squarks and is divided into 24 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$ and the number of jets; MB-GGd which targets production of gluinos and is divided into 18 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$; MB-C which targets compressed spectra and is divided into 18 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$ and the number of jets. The control regions are also implemented, which gives an opportunity to fully exploit the full likelihood model provided by the ATLAS Collaboration.
Figure 8 shows the validation plots for the squark pair-production process. The simplified shape-fit has excellent agreement with the official result, while the full likelihood model gives a somewhat weaker exclusion. In either case, there is a clearly visible advantage over the best-SR method.
@@
This is a search [38] for squarks and gluinos in final states with one isolated lepton, jets, and missing transverse momentum. Benchmark models assume long decay chains for squarks and gluinos with charginos, neutralinos and gauge bosons in intermediate states that give rise to the final state lepton. There is one multibin signal region that combines 26 bins defined according to the number of jets, number of identified $b$-jets, and $m_{\textrm{eff}}$. The full likelihood model is provided.
The validation plots for gluino pair production, followed by the decay to intermediate mass chargino are shown in Fig. 9. The shape of the ATLAS exclusion line is well reproduced for the likelihood models, whereas the BSR exclusion is significantly weaker.
@@
This is a search [39] for invisible decays of a Higgs boson or dark matter (DM) particles, $\chi$, produced in association with a $Z$ boson. The final state $Z$ boson is reconstructed from a same-flavor opposite-sign dilepton pair. One multibin signal region, optimized for the 2HDM+$a$ model [80], is implemented in CheckMATE. It is divided into 22 bins according to transverse mass, $m_{\mathrm{T}}$. The implementation is based on the simplified likelihood model.
@@
This is a search [40] for supersymmetric particles (squarks and gluinos) in final states with jets ($\geq 2$) and a large missing transverse momentum. The events are required to have at least one energetic jet. For the inclusive search, events are required to have at least 2 jets and signal regions are based on the number of jets, $b$-jets, $H_{T}$, and the $M_{T2}$ variable calculated using two pseudojets. Only the signal regions for prompt production are currently implemented in CheckMATE. There are 21 inclusive aggregate signal regions and 282 exclusive bins which are combined into two multi-bin signal regions using a covariance matrix. Due to numerical problems with the entire matrix $282\times 282$, we split the calculation into two multibin signal regions (corresponding to final states with low- and high-$H_{T}$, respectively) that are marginally correlated [81]. However, these two signal regions can be further combined by a user with an option SRCombination.
@@
For validation, we compare exclusion limits for the vector mediator model in the $m_{\textrm{mediator}}$$m_{\textrm{DM}}$ plane. The mediator-quarks and mediator-DM couplings are fixed to $g_{q}=0.25$ and $g_{\chi}=1$, respectively. The validation plot, Figure 13, again shows good agreement between the CheckMATE and CMS limits.
@@
This is a search [82] for charginos and neutralinos, that looks for final states with large missing transverse momentum and pairs of hadronically decaying bosons $WW$, WZWZ and $WH$. This search makes use of specific algorithms (taggers) defined to identify $W$ boson/Higgs boson candidates out of the identified large-radius signal jets. The individual bins are defined orthogonally based on the number of $b$-tagged jets, the number of identified $W$, $Z$, and Higgs boson candidates and missing transverse momentum. The simplified likelihood is built out of 35 signal regions. Additionally, there are 4 aggregate signal regions defined.
@@ 4 Conclusions and outlook
In this paper, we presented a version of CheckMATE, which significantly updates the database of available experimental searches. All included searches are based on the full data set of LHC Run 2. Here, we describe the implementation of 9 ATLAS searches and 4 CMS searches. Another update comes with the inclusion of statistical models, which in a number of examples significantly improve sensitivity and expand exclusion limits. This implementation includes full and simplified likelihoods provided by the ATLAS collaboration, and simplified correlated background models provided by CMS. With a number of options available, it allows users to fully control the numerical evaluation. CheckMATE can be downloaded from the GitHub repository:
@@
More information, including previous versions, expanded validation notes can be found at:
@@
In future releases, we plan to further expand CheckMATE’s capabilities by allowing statistical combination of different searches and orthogonal signal regions. In a separate note we will cover another new development to CheckMATE, which is an interface to machine learning methods which are becoming increasingly common tool in experimental analyses at the LHC.
@@ Acknowledgements
The authors thank Zirui Wang for his help and additional input for Ref. [39]. This work was partially supported by the National Science Centre, Poland under grant 2019/35/B/ST2/02008. The authors have also received funding from the Norwegian Financial Mechanism 2014-2021, grant 2019/34/H/ST2/00707.

### v3

@@ 3 Validation
In Figure 4, we show validation plots for slepton search signal regions. The electron and muon channels are shown separately. Very good agreement between ATLAS and CheckMATE is obtained for the shape-fit analysis. Additionally, by comparing red exclusion curves obtained using the best-SR-method, we find a clear benefit of using multibin signal regions. Since there is a negligible difference between expected and observed results (both for ATLAS and CheckMATE), we only show the observed limits.
In Figure 5, we show validation plots for the production of wino-like chargino-neutralino pairs (EW search region). The searches for different parities of LSP and NLSP are shown separately: upper row $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})>0$ (same parity) and $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})<0$ (opposite parity) in the lower row; see, e.g. Ref. [53] for a discussion of CP properties of the neutralino sector and consequences for the decay kinematics. Clearly, the simplified likelihood has a problem with reproducing the complicated shape of the observed limit due to significant excesses and deficits in several signal regions. Note that a similar conclusion was obtained by ATLAS in Ref. [74]. Therefore, for this search, we advise against using the simplified likelihood in routine runs. The agreement between the full likelihood and the ATLAS observed limit for the case of same parity neutralinos (top row) is not ideal as well, even though the expected limits comparison is rather encouraging and within 1-$\sigma$ uncertainty, see the bottom plot in Fig. 5. The search is also included in HackAnalysis, cf. Ref. [14], SModelS [19], and MadAnalysis [24]. Additional validation material for CheckMATE, including cut-flows, can be found in [93].
Figure 4: Validation plots (left: observed limits; right: expected limits) for the search atlas_1911_12606 (SUSY-2018-16), slepton pair production. Top row: smuon search; bottom row: selectron search. Left panel: observed limits; right panel: expected limits.
Figure 5: Validation plots for the search atlas_1911_12606 (SUSY-2018-16). Top row: $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})>0$; middle row: $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})<0$. Left panels: observed limits; right panels: expected limits. In the bottom row comparison of expected 1-$\sigma$ bands of ATLAS and CheckMATE full likelihood model for the same-parity case.
@@
This is a search [4] for hadronically decaying supersymmetric partners of top quark, and up-type, 3rd generation scalar leptoquark. The final states consist of several jets, 0 leptons, and large $E_{\mathrm{T}}^{\mathrm{miss}}$. The requirements for $b$ jets vary between signal regions. There are 3 multibin signal regions: SRA-B that targets scenarios with highly boosted top quarks in the final state and is divided into 6 bins according to an invariant mass of a large-$R$ jet; SRC, for scenarios with 3-body decays of stops, that is divided into 5 bins according to a recursive jigsaw reconstruction technique variable $R_{\textrm{ISR}}$ [75]; SRD for scenarios with compressed spectra and 4-body decays of stops which is divided into 3 bins according to the number of identified $b$-jets. No likelihood model is provided, and only simplified likelihood fitting is available. The combinations are performed separately for each signal category (corresponding to three likelihood models), A-B, C, and D, as described in the ATLAS paper. This avoids overlaps and accidental correlations.
The validation plots for this search are shown in Figure 6. We include the full range of stop masses and decay modes. Good agreement across the parameter plane is observed. For the high-stop-mass region the CheckMATE result is slightly weaker than ATLAS, clearly extending the exclusion reach compared to single SR exclusion. In more compressed scenarios, the simplified fit gives a slightly too strong exclusion in some regions of the parameter space; however, this is below 30% difference in the exclusion strength. Additional validation material can be found in [92].
@@
[…]
@@
Figure 7: Validation plots for the search atlas_2006_05880 (SUSY-2018-21). Left panels: observed limits; right panels: expected limits.
@@
This is a search [9] for squarks (1st and 2nd generation) and gluinos in final states with 2-6 jets, 0 leptons, and missing transverse momentum. There are three multibin signal regions in this search: MB-SSd which targets squark pair production and is divided into 24 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$ and the number of jets; MB-GGd which targets gluino production and is divided into 18 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$; MB-C which targets compressed spectra and is divided into 18 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$ and the number of jets. The control regions are also implemented, which gives an opportunity to fully exploit the full likelihood model provided by the ATLAS Collaboration.
Figure 8 shows the validation plots for the squark pair-production process. The simplified shape-fit has excellent agreement with the official result, while the full likelihood model gives a somewhat weaker exclusion. In either case, there is a clearly visible advantage over the best-SR method. In Appendix we provide detailed signal yields for the multibin SRs of this search. Additional validation material (cut flows) can be found in [94].
@@
This is a search [10] for squarks and gluinos in final states with an isolated lepton, jets, and missing transverse momentum. Benchmark models assume long decay chains for squarks and gluinos with charginos, neutralinos and gauge bosons in intermediate states that give rise to the final state lepton. There is a multibin signal region that combines 26 bins defined according to the number of jets, the number of identified $b$-jets, and $m_{\textrm{eff}}$. The full likelihood model is provided.
The validation plots for gluino pair production, followed by the decay to intermediate mass chargino, are shown in Fig. 9. The shape of the ATLAS exclusion line is well reproduced for the likelihood models, whereas the BSR exclusion is significantly weaker.
@@
This is a search [11] for invisible decays of a Higgs boson or dark matter (DM) particles, $\chi$, produced in association with a $Z$ boson. The final state $Z$ boson is reconstructed from a same-flavor opposite-sign dilepton pair. A multibin signal region, optimized for the 2HDM+$a$ model [33], is implemented in CheckMATE. It is divided into 22 bins according to the transverse mass, $m_{\mathrm{T}}$. The implementation is based on the simplified likelihood model.
@@
This is a search [54] for supersymmetric particles (squarks and gluinos) in final states with jets ($\geq 2$) and a large missing transverse momentum. The events are required to have at least one energetic jet. For the inclusive search, events are required to have at least 2 jets and signal regions are based on the number of jets, $b$-jets, $H_{T}$, and the $M_{T2}$ variable calculated using two pseudojets. Only the signal regions for prompt production are currently implemented in CheckMATE. There are 21 inclusive aggregate signal regions and 282 exclusive bins, which are combined into two multi-bin signal regions using a covariance matrix. Due to numerical problems with the entire matrix $282\times 282$, we split the calculation into two multibin signal regions (corresponding to final states with low and high-$H_{T}$, respectively) that are marginally correlated [85]. However, these two signal regions can be further combined by a user with the SRCombination input parameter, and the UnCorrStatisticsCombiner function of the Spey package.
@@
For validation, we compare exclusion limits for the vector mediator model in the $m_{\textrm{mediator}}$$m_{\textrm{DM}}$ plane. The mediator-quark and mediator-DM couplings are fixed to $g_{q}=0.25$ and $g_{\chi}=1$, respectively. The validation plot, Figure 13, again shows good agreement between the CheckMATE and CMS limits.
@@
This is a search [98] for charginos and neutralinos that looks for final states with large missing transverse momentum and pairs of hadronically decaying bosons $WW$, WZWZ and $WH$. This search makes use of specific algorithms (taggers) defined to identify $W$ boson/Higgs boson candidates out of the identified large-radius signal jets. Individual bins are defined orthogonally based on the number of $b$-tagged jets, the number of identified $W$, $Z$, and Higgs boson candidates, and missing transverse momentum. The simplified likelihood is built out of 35 signal regions. Additionally, there are 4 aggregate signal regions defined.
@@ 4 Conclusions and outlook
In this paper, we present a version of CheckMATE, which significantly updates the database of available experimental searches. All included searches are based on the full data set of LHC Run 2. Here, we describe the implementation of 9 ATLAS searches and 4 CMS searches. Another update comes with the inclusion of statistical models, which in a number of examples significantly improve sensitivity and expand exclusion limits. This implementation includes the full and simplified likelihoods provided by the ATLAS Collaboration and the simplified correlated background models provided by CMS. With a number of options available, it allows users to fully control the numerical evaluation. CheckMATE can be downloaded from the GitHub repository:
@@
More information, including previous versions and expanded validation notes, can be found at:
@@
In future releases, we plan to further expand the CheckMATEcapabilities by allowing a statistical combination of different searches and orthogonal signal regions. In a separate note, we will cover another new development to CheckMATE, which is an interface to machine learning methods which are becoming increasingly common tools in experimental analyses at the LHC.
@@ Acknowledgements
The authors thank Zirui Wang for his help and additional input for Ref. [11]. This work was supported by the National Science Centre, Poland under grant 2019/35/B/ST2/02008 and OpenMAPP project under CHIST-ERA programme (grant No. NCN 2022/04/Y/ST2/00186). The authors have also received funding from the Norwegian Financial Mechanism 2014-2021, grant 2019/34/H/ST2/00707.
This version of the article has been accepted for publication after peer review, but is not the Version of Record and does not reflect post-acceptance improvements, or any corrections. The Version of Record is available online at: https://doi.org/10.1140/epjc/s10052-026-15414-8.
Appendix
Event yields for atlas_2010_14293 (SUSY-2018-22)
In Tables 68 we provide a detailed comparison of the event yields in the signal regions of the atlas_2010_14293 (SUSY-2018-22) search where the MC sample size in CheckMATE was 100k. A good agreement (within statistical uncertainties) between ATLAS and CheckMATE is observed except for the 4-jet bins in the squark search region MB-SSd. However, the test sample, the squark pair production with direct decays $\tilde{q}\to q\tilde{\chi}_{1}^{0}$, at the lowest parton order does not contribute to these bins and requires 2 additional jets from ISR/FSR. It is therefore very sensitive to details in MC modeling and is difficult to reproduce in recasts. These discrepancies do not affect the fit for this signal process due to the small yields compared to expected backgrounds.
Bin CheckMATE ATLAS ATLAS/CheckMATE
MB-SSd-2-1000-10 $0.94\pm 0.033$ $\pm 0.15 1.1$$1\pm 0.16$
MB-SSd-2-1000-16 $0.74\pm 0.029$ $0.59\pm 0.1$ $0.8\pm 0.14$
MB-SSd-2-1000-22 $0.41\pm 0.022$ $0.27\pm 0.071$ $0.66\pm 0.17$
MB-SSd-2-1600-10 $3.8\pm 0.066$ $3.7\pm 0.26$ $0.96\pm 0.069$
MB-SSd-2-1600-16 $2.2\pm 0.05$ $2\pm 0.18$ $0.91\pm 0.085$
MB-SSd-2-1600-22 $4\pm 0.067$ $3.5\pm 0.26$ $0.88\pm 0.064$
MB-SSd-2-2200-16 $1.8\pm 0.045$ $2\pm 0.2$ $1.1\pm 0.11$
MB-SSd-2-2200-22 $7.8\pm 0.094$ $8.4\pm 0.39$ $1.1\pm 0.051$
MB-SSd-2-2800-16 $1.4\pm 0.04$ $1.3\pm 0.16$ $0.95\pm 0.11$
MB-SSd-2-2800-22 $5.9\pm 0.082$ $6\pm 0.33$ $1\pm 0.055$
MB-SSd-2-3400-22 $1.2\pm 0.038$ $0.93\pm 0.14$ $0.77\pm 0.12$
MB-SSd-2-3400-28 $6\pm 0.055 2.2$$2\pm 0.19$ $0.84\pm 0.075$
MB-SSd-2-4000-22 $0.47\pm 0.024$ $0.22\pm 0.063$ $0.47\pm 0.13$
MB-SSd-2-4000-28 $0.89\pm 0.033$ $0.67\pm 0.11$ $0.76\pm 0.12$
MB-SSd-4-1000-10 $0.13\pm 0.012$ $0.63\pm 0.1$ $4.7\pm 0.78$
MB-SSd-4-1000-16 $0.067\pm 0.0088$ $0.39\pm 0.08$ $5.8\pm 1.2$
MB-SSd-4-1000-22 $0.015\pm 0.0041$ $0.17\pm 0.064$ $11\pm 4.3$
MB-SSd-4-1600-10 $1.7\pm 0.045$ $2.4\pm 0.21$ $1.4\pm 0.13$
MB-SSd-4-1600-16 $0.52\pm 0.025$ $1.3\pm 0.16$ $2.6\pm 0.3$
MB-SSd-4-1600-22 $0.39\pm 0.021$ $1.7\pm 0.17$ $4.3\pm 0.44$
MB-SSd-4-2200-16 $0.77\pm 0.03$ $1.4\pm 0.16$ $1.8\pm 0.21$
MB-SSd-4-2200-22 $1.6\pm 0.042$ $3.9\pm 0.26$ $2.5\pm 0.17$
MB-SSd-4-2800-16 $1.1\pm 0.036$ $0.99\pm 0.14$ $0.92\pm 0.13$
MB-SSd-4-2800-22 $1.9\pm 0.047$ $4.6\pm 0.3$ $2.4\pm 0.16$
Table 6: Event yields for the squark pair production, $m_{\tilde{q}}=1700$ GeV, $m_{\tilde{\chi}_{1}^{0}}=500$ GeV with MC statistical uncertainties, $s\pm\Delta s$, normalized to the nominal cross section (8-fold degeneracy) at the approximate NNLO+NNLL accuracy, $\sigma=0.74$ fb.
Bin CheckMATE ATLAS ATLAS/CheckMATE
MB-GGd-4-1000-10 $4\pm 0.13$ $2.9\pm 0.49$ $0.74\pm 0.12$
MB-GGd-4-1000-16 $1.4\pm 0.076$ $0.64\pm 0.21$ $0.46\pm 0.15$
MB-GGd-4-1000-22 $18\pm 0.027 0.0$$097\pm 0.097$ $0.55\pm 0.55$
MB-GGd-4-1600-10 $17\pm 0.26$ $17\pm 1.3$ $1\pm 0.076$
MB-GGd-4-1600-16 $13\pm 0.23$ $4\pm 1.1 1.1$$1\pm 0.086$
MB-GGd-4-1600-22 $5.1\pm 0.15$ $4.4\pm 0.63$ $0.85\pm 0.12$
MB-GGd-4-2200-10 $11\pm 0.21$ $13\pm 1.1$ $1.2\pm 0.1$
MB-GGd-4-2200-16 $14\pm 0.24$ $17\pm 1.2$ $1.2\pm 0.088$
MB-GGd-4-2200-22 $12\pm 0.23$ $4\pm 1.1 1.1$$1\pm 0.086$
MB-GGd-4-2800-10 $2.6\pm 0.11$ $2.4\pm 0.42$ $0.91\pm 0.16$
MB-GGd-4-2800-16 $3.6\pm 0.12$ $3.6\pm 0.54$ $1\pm 0.15$
MB-GGd-4-2800-22 $4.3\pm 0.13$ $5.3\pm 0.65$ $1.2\pm 0.15$
MB-GGd-4-3400-10 $0.49\pm 0.046$ $0.56\pm 0.21$ $1.1\pm 0.42$
MB-GGd-4-3400-16 $0.47\pm 0.046$ $0.3\pm 0.14$ $0.64\pm 0.3$
MB-GGd-4-3400-22 $0.71\pm 0.056$ $1.5\pm 0.36$ $2.1\pm 0.51$
MB-GGd-4-4000-10 $0.11\pm 0.023$ $0.24\pm 0.14$ $2.2\pm 1.3$
MB-GGd-4-4000-16 $1\pm 0.022 0.11±0$$1\pm 0.1$ $0.98\pm 0.95$
MB-GGd-4-4000-22 $0.21\pm 0.031$ $0.1\pm 0.1$ $0.48\pm 0.48$
Table 7: Event yields for the gluino pair production, $m_{\tilde{g}}=1800$ GeV, $m_{\tilde{\chi}_{1}^{0}}=800$ GeV with MC statistical uncertainties, $s\pm\Delta s$, normalized to the nominal cross section at the approximate NNLO+NNLL accuracy, $\sigma=2.93$ fb.
Bin CheckMATE ATLAS ATLAS/CheckMATE
MB-C-2-1600-16 $2.2\pm 0.05$ $1.9\pm 0.18$ $0.89\pm 0.084$
MB-C-2-1600-22 $9\pm 0.066 3.3$$3\pm 0.24$ $0.85\pm 0.062$
MB-C-2-2200-16 $2.1\pm 0.049$ $2.3\pm 0.22$ $1.1\pm 0.1$
MB-C-2-2200-22 $9.1\pm 0.1$ $9.8\pm 0.42$ $1.1\pm 0.047$
MB-C-2-2800-16 $1.7\pm 0.044$ $1.5\pm 0.17$ $0.91\pm 0.098$
MB-C-2-2800-22 $9.5\pm 0.1$ $9.5\pm 0.41$ $1\pm 0.044$
MB-C-4-1600-16 $0.76\pm 0.029$ $0.71\pm 0.12$ $0.93\pm 0.16$
MB-C-4-1600-22 $0.69\pm 0.028$ $0.63\pm 0.11$ $0.91\pm 0.16$
MB-C-4-2200-16 $\pm 0.034 1.1$$\pm 0.14 1.1$$1\pm 0.14$
MB-C-4-2200-22 $2.9\pm 0.058$ $2.5\pm 0.21$ $0.86\pm 0.07$
MB-C-4-2800-16 $1\pm 0.035$ $0.76\pm 0.13$ $0.73\pm 0.12$
MB-C-4-2800-22 $4.7\pm 0.074$ $4\pm 0.28$ $0.84\pm 0.06$
MB-C-5-1600-16 $0.35\pm 0.02$ $0.28\pm 0.07$ $0.8\pm 0.2$
MB-C-5-1600-22 $0.19\pm 0.015$ $0.24\pm 0.068$ $1.2\pm 0.35$
MB-C-5-2200-16 $1\pm 0.034$ $0.61\pm 0.11$ $0.61\pm 0.11$
MB-C-5-2200-22 $1.6\pm 0.043$ $1.2\pm 0.15$ $0.77\pm 0.094$
MB-C-5-2800-16 $1.5\pm 0.043$ $0.75\pm 0.13$ $0.5\pm 0.085$
MB-C-5-2800-22 $4.6\pm 0.074$ $3.3\pm 0.25$ $0.71\pm 0.055$
Table 8: Event yields for the squark pair production, $m_{\tilde{q}}=1700$ GeV, $m_{\tilde{\chi}_{1}^{0}}=500$ GeV with MC statistical uncertainties, $s\pm\Delta s$, normalized to the nominal cross section (8-fold degeneracy) at the approximate NNLO+NNLL accuracy, $\sigma=0.74$ fb.